OpenStudy (anonymous):

Help...1. Select two ordered pair points of your own. (Just pick any two points that seem relatively easy to work with.) Be sure to list these points for your instructor. Find the slope and show your work.

OpenStudy (anonymous):

Just need someone to help me pick points and then I will solve the slope and then ask you if I got it right

OpenStudy (solomonzelman):

Slope formula is \[\huge\color{red}{\frac{y_1-y_2}{x_1-x_2}=m}~~~~~(m~~is~~the~~slope)\] Lets choose (2,4) AND (1,-2)

OpenStudy (solomonzelman):

\[\huge\color{red}{\frac{4-(-2)}{2-1}}~~~~->~~~~\huge\color{red}{\frac{6}{1}}\] So it would be equal to 6.

OpenStudy (anonymous):

Okay so the slope is 6 or 6/1

OpenStudy (solomonzelman):

Yes, do you see how I plugged in the points?

OpenStudy (anonymous):

yeah you put both of the y's on the top and subtracted them and both the x's on the bottom and subtrcted them

OpenStudy (anonymous):

subtracted*

OpenStudy (solomonzelman):

Yes, CORRECT!

OpenStudy (solomonzelman):

Do you want to learn the rest about these equations?

OpenStudy (anonymous):

sure what do u got

OpenStudy (solomonzelman):

OK, lets first talk about the y-intercept.

OpenStudy (anonymous):

kay

OpenStudy (solomonzelman):

y- intercept is the point on the line where it intersects the y-axis. For example, 

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

If you want to find and lets say you where given the slope but and one point through which it passes.

OpenStudy (solomonzelman):

(The slope-intercept form) y=mx+b m= slope (we already know that) b= y-intercept So if we have \[y=4x+b\] and it passes through the point (2,1), to find the y-intercept you would plug in the point.

OpenStudy (anonymous):

so 1 would go where the b is ?

OpenStudy (solomonzelman):

I got disconnected, sorry.

OpenStudy (anonymous):

Its fine

OpenStudy (solomonzelman):

Look at (x,y) and compare it to (2,1) to plug it into your equation. \[y=4x+b~~~and~~~i t~~~goes~~~through~~~\color{red} {(2,1)}\]2 is the x coordinate 1 is the y coordinate \[y=4x+b~~~~~->~~~~~\color{red} {1}=4(\color{red} {2})+b\]\[1=8+b\]\[1\color{red} {-8}=8+b\color{red} {-8}\]\[\color{blue} {7=b}\] knowing that b=7 we can put it into \[\color{black} {y=4x+b}~~~~~->~~~~~y=4x+\color{blue} {7}\]

OpenStudy (solomonzelman):

See how I fond the y-intercept (tell me after you go over the explanation)